Status Report

NASA Space Station On-Orbit Status 3 Jan 2004

By SpaceRef Editor
January 3, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 3 Jan 2004

All ISS systems continue to function nominally, except as noted previously or below.

CDR/SO Michael Foale and FE Alexander Kaleri performed the regular 3-hr. Saturday task of station cleaning.  [“Uborka stantsii” focuses on removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with disinfectants (“Fungistat”) and cleaning of fan screens to avoid temperature rises.]

Procedures were uplinked for Kaleri for manual restart of the Elektron-VM oxygen generator from its controlling laptop, in case of anotherotkaza(shutdown).

If required for maintaining the proper ppO2 level, today’s schedule also provided for burning two more SFOG (solid-fuel oxygen generator) candles.

Working off the Russian discretionary task list, Sasha performed another set of observations for the Diatomeya ocean research program, taking imagery depicting bioproductive waters of the Atlantic and Indian oceans and of the hydro-meteorological background.  [The FE focused the Nikon F5 with f/80 mm lens from SM window #8 and the Sony DVCAM-150 digital camcorder, fixed on a bracket above window #7, on the island province in the western part of the Indian Ocean along the axis of the Tanzania shelf plate, the Comoro Islands and the littoral water areas of Madagascar.  Other targets were the high-pressure system vortex in the Indian Ocean off the coast of South Africa towards the Kerguelen Island, in the Atlantic Ocean the north-western coast of Africa from the Capo Verde islands to the equator, as well as the Falklands-Patagonia bio-productive region.]

As a second task list item, Kaleri performed another round of Uragan earth observations, using the Kodak 760 electronic still camera (ESC) with the long f800 lens at window #9 to make observations and take imagery of the earthquake region of the Iranian city of Bam.  [Because of the destruction, Bam now appears only as a large number of dark spots on the yellow/red landscape background on the banks of a small river on the intersection of two highways and several local roads.  Other Uragan target areas today were such locations as Tunisia, Libya, Chad, Sudan, Uganda, Madagascar, the Bahamas, Brazil, Guatemala, Ecuador, Bolivia, Chile and Argentina.]
The FE had his regular weekly PFC (private family conference), via Russian VHF.

Kaleri also conducted the daily routine maintenance of the SOZh life support system and ASU toilet facilities.

The crew conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

As a long-term item on the Russian task list, Kaleri is charged with taking photographs of the PKZ-1V Kromka experiment tablet deployed on the plume deflector of the SM’s plus-pitch thrusters, as was Yuri Malenchenko before him. [The pictures are taken with the Kodak 760 digital still camera (DSC) from the EVA hatch 1 window in the DC-1 docking compartment and downlinked via OCA.  Kromka collects contamination effluents from the thrusters on a plate for later analysis.]
During yesterday’s “ghosting” (cloning) of PCS HDD (portable computer system/hard disk drive) #6063 the drive performed extremely slow, and prior to sleep the crew reinstalled HDD #6054.  [According to Mike’s report this morning, two attempts to ghost the drive using an auto loader floppy failed.  The crew believes that the floppy has a problem and recommends using a different one for future ghosting.  Because of its very slow loading, there is probably also a hardware problem with the HDD.  Specialists are to develop a troubleshooting plan for both issues.]

Component swapping and resizing for the U.S. EMUs (extravehicular mobility units) is scheduled for 1/5 (Monday), in order to have the spacesuits ready in case of any unscheduled EVA.  It is left up to the crew to perform this task already over the weekend.

In the SM, TsUP/Moscow early this morning commanded battery #1 to cycle mode.

Both crewmembers worked out with their daily 2.5-h program of physical exercise, on TVIS treadmill, RED expander, and VELO cycle with load trainer.

Weekly Science Update (Expedition Eight — 9th):

GASMAP:   Next scheduled 30-day health check will be coming up soon.

Human Research Facility/Workstation (HRF WS):  Continuing.

Advanced Ultrasound:  Planned.

Hand Posture Analyzer (HPA): Looking forward to the new sessions this year.

In-Space Soldering Investigation (ISSE): Thanks to Mike for completing the ISSI Coupon Solder Wrap.  The ground team appreciated being “let in” into the Lab while Foale as working, and the digital pictures taken by him are “great”.  Also: “the Coupons all look wonderful and we are looking forward to starting the actual soldering operations”.

Pore Formation and Mobility Investigation (PFMI):  Thanks to Mike Foale for the successful repair of the PFMI Thermal Chamber Gradient Stage primary gear.  The team is looking forward to the next PFMI run next week.
Space Acceleration Measurement System (SAMS):   Mike was thanked for his hard work and troubleshooting in support of SAMS effort of characterizing the microgravity environment.
Microgravity Acceleration Measurement System (MAMS):  The MAMS filter cleaning went smoothly and MAMS is back in nominal data capturing mode.    

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   Planned.

Renal Stone (RS):  Next in-flight RS operations are scheduled for the weekend of 1/9-1/12.  Per crew request, the hardware set-up time on 1/9 was extended to an hour to allow performance of a hardware inventory.  The RS team will be supporting this activity if any questions or comments arise.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):  Completed.

Foot/Ground Reaction Forces during Space Flight (FOOT):  Good news from Mike’s first on-orbit FOOT session: He collected just over 6.5 hours of data with very few anomalies.  There was no sign of data loss due to electrode loosening and all channels gave good data.  “All the button presses were in the right places!”

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): The CBOSS-FDI team and the NIH (National Institute of Health) PI sincerely appreciate all of the crew’s efforts on FDI this week.  They are still in the process of analyzing the video and images that the crew took throughout this week and last.  Once a complete analysis of the data has been completed, the PI would like to give Mike a short summary of what has been accomplished so far.  There will be no FDI activities in the coming two weeks, as the data analysis completion is being awaited.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Students are looking forward to the next session later this Increment.

Earth Science Toward Exploration Research (ESTER):  Looking forward to the sessions next year.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Completed.

Miscible Fluids in Microgravity (MFMG):    Sasha was thanked for heating up the honey that will be used in the MFMG Experiment  by Mike in the near future.

Educational Payload Operations (EPO):  The Educators are excited about using the video from the crew’s education demonstrations in classrooms and workshops.  They were thanked for providing this valuable education resource.  The ground is looking forward to upcoming activities.
For today’s CEO (crew earth observations) sessions, the ground uplinked a list of target cities as potential photographic opportunities at the crew’s discretion.

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 6:49am EST [= epoch]):

  • Mean altitude — 369.4 km
  • Apogee — 373.9 km
  • Perigee — 364.9 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006658
  • Orbits per 24-hr. day — 15.66
  • Mean altitude loss in last 24 hours — 135
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29232

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.